Q6. A system composed of 2 kg of the above fluid expands in a frictionless piston and cylinder 

machine from an initial state of 1 MPa, 100°C to a final temperature of 30°C. If there is no heat 

transfer, find the net-work for the process.


1
Expert's answer
2021-01-22T08:11:07-0500

"pV=\frac{m}{M}RT"

"\frac {p_1V_1}{T_1}=\frac {p_2V_2}{T2}"


"p_1=1 MPa"


"t_1=100^0C"


"t_2=30^0C"

"p_2 \frac {V_2}{V_1}=\frac {p_1T_1}{T_2}=1.231 MPa"